# Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis

### v1

@@ Conclusion
We introduced Retro-Expert, the first retrosynthesis framework to address long-standing interpretability limitations by generating human-readable, step-by-step reasoning. Our approach synergizes specialized models, which construct a chemical decision space, with an LLM that navigates it using a policy trained via RL. Experiments show Retro-Expert achieves competitive accuracy and strong generalization while allowing for plug-and-play modularity. By bridging the gap between AI prediction and a chemist’s workflow, Retro-Expert establishes a new paradigm for trustworthy and collaborative AI in chemical discovery.
Acknowledgments
This work was supported by the National Science and Technology Major Project (2023ZD0120802).
Appendix
@@ Acknowledgments
This supplemental document provides details on our Retro-Expert. In Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis, we describe the technical implementation of Retro-Expert. Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis presents all the prompts used in LLM-based baselines. In Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis, we visualize representative reasoning examples to analyze the reasoning capabilities of Retro-Expert, including successful selection and generation cases, as well as failure cases. Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis details the interpretability evaluation setup, covering the prompts corresponding to the evaluation metrics, and human evaluation procedures. In Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis, we provide the prompts applied for all models when evaluated on the out-of-distribution (OOD) dataset. Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis shows the details of our web lab experiments. Finally, Section Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis presents additional ablation studies.
[…]
A. Details of KGPO Module
This section elaborates on the technical implementation details of our Retro-Expert.
@@ A. Details of KGPO Module
Training Samples. To optimize training efficiency and enhance the model’s ability to learn diverse chemical reaction principles, we did not utilize the entire training dataset directly. Instead, we meticulously curated a high-quality training subset. The original dataset exhibited a significant long-tailed distribution of reaction templates and severe imbalance in sample quantities across different reaction types.
@@
Training Configuration. We employ reinforcement learning via GPRO (Generalized Reinforcement Learning with Policy Optimization) to optimize the reasoning strategy of the LLM (Qwen-7B-Instruct). To ensure training stability while promoting deep deliberation and exploration, the core hyperparameters are configured as follows: train_epochs=2, max_completion_length=8192, temperature=0.7, $\beta$=0.001.
@@
The prompt used for training is as follows:
System: You are a helpful AI Assistant that provides well-reasoned and detailed responses. You first think about the reasoning process as an internal monologue and then provide the user with the answer. Respond in the following format: $\n...\n</think>\n$$\n...\n</answer>$ ...................................................................................................................................................... Assistant: You are an experienced chemist analyzing chemical retrosynthesis. Given the standard SMILES representation of the product is: {Product Standard SMILES}, Its automapping version is: {Product Mapped SMILES}, The IUPAC name of the product is: {IUPAC Name}. The most likely reaction type for synthesizing this product is: {Reaction Type}. The most likely reaction center for synthesizing this product is: {Reaction Center}. Reactants Candidate Set:
@@
1. The prompt for our Retro-Expert training.
[…]
B. Details of Comparison Methods
This section details all prompts employed in the LLM-based methods within our comparative experiments. Specifically, for chemical LLMs, which inherently adhere to prescribed dialogue schemas. To ensure valid predictions, we strictly follow each model’s native evaluation templates when generating output.
2. The prompt for nach0 and ChemCrow.
@@ B. Details of Comparison Methods
4. The prompt for ChemDFM.
You are an experienced chemist analyzing chemical retrosynthesis. Given the standard SMILES representation of the product is: {Product Standard SMILES}, Its automapping version is: {Product Mapped SMILES}, The IUPAC name of the product is: {IUPAC Name}. The most likely reaction type for synthesizing this product is: {Reaction Type}. The most likely reaction center for synthesizing this product is: {Reaction Center}. Reactants Candidate Set:
@@
5. The prompt for general LLM.
[…]
C. Results Analysis
@@
[…]
C.1. Successful Reasoning (Selection Scene)
We demonstrate the reasoning processes generated by our Retro-Expert across diverse reaction types (in Figure 4-9). This indicates Retro-Expert’s fundamental competency in executing chemically valid reasoning and making accurate selections within a high-quality chemical action space.
@@ C.1. Successful Reasoning (Selection Scene)
[…]
C.2. Successful Reasoning (Generation Scene)
Furthermore, we showcase Retro-Expert’s capacity for self-reflection and generation. Crucially, when confronted with suboptimal candidates from all specialized models, Retro-Expert engages in critical analysis and autonomously synthesizes a novel, chemically sound solution (in Figure 10-12).
@@ C.3. Failure Analysis
(b) Failure case for multiple possible reaction paths.
@@
[…]
D. Details of Interpretability Evaluation
@@
[…]
D.1. Details of Evaluation Metrics
@@
Mechanism Accuracy ($\mathbb{MA}$) System: You are an expert in chemical retrosynthetic analysis, responsible for evaluating the quality of reasoning processes generated by retrosynthetic models. #INSTRUCTIONS: Based on the following the Generated Reasoning Process, conduct a professional assessment across the follow dimensions: Mechanism Accuracy: Assess whether the reaction mechanism described in the reasoning aligns with established chemical principles. [Score 1]: Incorrect mechanism. [Score 2]: Partially correct but with significant deviations. [Score 3]: Partially correct with minor deviations. [Score 4]: Largely consistent. [Score 5]: Fully consistent. Evaluations must be grounded in given Authentic Chemical Synthesis Information and the results must be strictly formatted according to the specified Output Format. [Output Format Example] Mechanism Accuracy:2 ........................................................................................................................................................... Assistant: Please evaluate the mechanism accuracy of the Generated Reasoning Process, based on the given Authentic Chemical Synthesis Information. {Generated Reasoning Process} Authentic Chemical Synthesis Information: The product’s standard SMILES representation is {Product Standard SMILES}; Its automapping version is {Product Mapped SMILES}; The IUPAC name of the product is {IUPAC Name}. The reaction type for synthesizing this product is: {The Golden Reaction Type}. The reaction center of the product molecule is: {The Golden Reaction Center}. The reactants for synthesizing the product molecule are: {The Golden Reactants}.
6. The prompt for Mechanism Accuracy ($\mathbb{MA}$).
@@
Factual Correctness ($\mathbb{FC}$) System: You are an expert in chemical retrosynthetic analysis, responsible for evaluating the quality of reasoning processes generated by retrosynthetic models. #INSTRUCTIONS: Based on the following the Generated Reasoning Process, conduct a professional assessment across the follow dimensions: Factual Correctness: Evaluate the presence of factual errors concerning chemical principles within the reasoning description. [Score 1]: Major factual errors. [Score 2]: Several errors. [Score 3]: minor error(s). [Score 4]: Essentially free of errors. [Score 5]: Completely accurate. Evaluations must be grounded in given Authentic Chemical Synthesis Information and the results must be strictly formatted according to the specified Output Format. [Output Format Example] Factual Correctness:2 ........................................................................................................................................................... Assistant: Please evaluate the factual correctness of the Generated Reasoning Process, based on the given Authentic Chemical Synthesis Information. {Generated Reasoning Process} Authentic Chemical Synthesis Information: The product’s standard SMILES representation is {Product Standard SMILES}; Its automapping version is {Product Mapped SMILES}; The IUPAC name of the product is {IUPAC Name}. The reaction type for synthesizing this product is: {The Golden Reaction Type}. The reaction center of the product molecule is: {The Golden Reaction Center}. The reactants for synthesizing the product molecule are: {The Golden Reactants}.
7. The prompt for Factual Correctness ($\mathbb{FC}$).
@@
Logical Consistency ($\mathbb{LC}$) System: You are an expert in chemical retrosynthetic analysis, responsible for evaluating the quality of reasoning processes generated by retrosynthetic models. #INSTRUCTIONS: Based on the following the Generated Reasoning Process, conduct a professional assessment across the follow dimensions: Logical Consistency: Determine if conclusions follow logically from premises, analyses, or assumptions without contradictory leaps. [Score 1]: Conclusion contradicts premises/analysis. [Score 2]: Partially consistent but with significant logical leaps. [Score 3]: Partially consistent with only minor leaps. [Score 4]: Largely fluent and natural. [Score 5]: Completely fluent and natural. Evaluations must be grounded in given Authentic Chemical Synthesis Information and the results must be strictly formatted according to the specified Output Format. [Output Format Example] Logical Consistency:2 ........................................................................................................................................................... Assistant: Please evaluate the logical consistency of the Generated Reasoning Process, based on the given Authentic Chemical Synthesis Information. {Generated Reasoning Process} Authentic Chemical Synthesis Information: The product’s standard SMILES representation is {Product Standard SMILES}; Its automapping version is {Product Mapped SMILES}; The IUPAC name of the product is {IUPAC Name}. The reaction type for synthesizing this product is: {The Golden Reaction Type}. The reaction center of the product molecule is: {The Golden Reaction Center}. The reactants for synthesizing the product molecule are: {The Golden Reactants}.
8. The prompt for Logical Consistency ($\mathbb{LC}$).
[…]
D.2. Details of Human Study
@@
[…]
E. Details of Out-of-Distribution Analysis
@@
9. The prompt for Out-of-Distribution Analysis.
[…]
F. Details of Web Lab Experiments
[…]
F.1. Synthesis of 3-(2-ethoxyphenyl)thiophene
Figure 15: The synthesis path of 3-(2-ethoxyphenyl)thiophene.
(a) 1H NMR spectrum
(b) 13C NMR spectrum
[…]
F.2. Synthesis of 1-(4-ethoxyphenyl)ethanone
Figure 17: The synthesis path of 1-(4-ethoxyphenyl)ethanone.
[…]
G. Additional Ablation Study
This section presents additional ablation studies.
Reward hacking. Due to the probabilistic bias in the Top-K candidate results provided by small models, i.e., the correct answer appears in the Top-1 position with significantly higher probability than in other positions. When using result correctness as the reward, the LLM will directly select the first candidate to continuously maximize its reward. To validate the effectiveness of our candidate position shuffling strategy, we visualized the positional distribution of correct answers in the small model’s K candidates and the distribution of positions selected by the LLM. As shown in Figure 19, using the original K candidates from small models results in severe reward hacking, whereas our position shuffling strategy effectively mitigates this issue.

### v2

@@ 5 Conclusion
We introduced Retro-Expert, a novel retrosynthesis framework to address long-standing interpretability limitations by generating human-readable reasoning process. Our approach synergizes specialized models, which construct a chemical decision space, with an LLM that navigates it using KGPO policy trained via RL. Experiments show Retro-Expert achieves competitive accuracy and strong generalization while allowing for plug-and-play modularity. By bridging the gap between AI prediction and a chemist’s workflow, Retro-Expert represents a significant step toward trustworthy and collaborative AI in chemical discovery.
Supplementary Material
@@
This supplemental document provides details on our Retro-Expert. In Section 6, we describe the technical implementation of Retro-Expert. Section 7 reports the performance of LLM-based baselines with and without the decision space, along with the prompts used for each baseline. In Section 8, we provide the details of our web lab experiments. Section 9 presents additional ablation studies, including the mitigation of reward hacking and the impact of varying N candidates on prediction performance. In Section 10, we visualize representative reasoning examples to analyze the reasoning capabilities of Retro-Expert, including successful selection and generation cases, as well as failure cases. Section 11 details the interpretability evaluation setup, covering the prompts corresponding to the evaluation metrics, and human evaluation procedures. Finally, Section 12, we provide the prompts applied for all models when evaluated on the out-of-distribution (OOD) dataset.
6 Details of KGPO Module
This section elaborates on the technical implementation details of our Retro-Expert. As depicted in Figure 5, we illustrate the optimization workflow of the KGPO module. This design underscores that the core of our KGPO is to promote critical thinking and reasoning grounded in chemical logic, rather than merely rewarding superficial correctness.
@@ 6 Details of KGPO Module
Training Samples. To optimize training efficiency and enhance the model’s ability to learn diverse chemical reaction principles, we did not utilize the entire training dataset directly. Instead, we meticulously curated a high-quality training subset. The original USPTO-50K dataset exhibited a significant long-tailed distribution of reaction templates and severe imbalance in sample quantities across different reaction types.
@@
For the USPTO-FULL dataset, we applied a similar strategy to obtain a training subset. Due to the absence of ground-truth reaction type labels, we conducted stratified random sampling solely based on the frequency of extracted reaction patterns, resulting in a final training set of 80k samples.
Training Configuration. We employ reinforcement learning via GPRO (Generalized Reinforcement Learning with Policy Optimization) to optimize the reasoning strategy of the LLM (Qwen2.5-7B-Instruct). To ensure training stability while promoting deep deliberation and exploration, the core hyperparameters are configured as follows: train_epochs=2, max_completion_length=8192, temperature=0.7, $\beta$=0.001.
@@
7 Details of Comparison Methods
This section additionally reports the prediction performance of all LLM-based methods under different space settings, along with the prompts used for each method. For chemical LLMs, which inherently adhere to predefined predictive patterns, we strictly follow each model’s evaluation templates to ensure valid predictions. The space-related information was incorporated in accordance with the respective evaluation templates.
Figure 5: Visualization of the KGPO’s optimization pipeline. Our target is to generate the final answer with an interpretable reasoning pathway by reasoning within the decision space. Thus we further implement a multi-stage reward mechanism that the reasoning text. The mechanism guides the model in two ways: it encourages rejecting incorrect candidates (e.g., “alkylation reaction would not” receives a reward of 1) while promoting inference of the correct reaction type (e.g., “deprotection reaction” receives a reward of 1).
Model Space Setting Top-1 Acc (%) BLEU Dis Validity MACCS RDK Morgan
LLM-based chemical models
T5Chem 43.64 0.972 8.42 0.982 0.956 0.974 0.925
gray!50[3pt/4pt]black ChemFormer 27.30 0.769 14.77 0.952 0.782 0.690 0.647
gray!50[3pt/4pt]black nach0 30.52 0.976 11.91 0.990 0.947 0.977 0.905
14.60 0.962 16.10 0.981 0.891 0.942 0.805
gray!50[3pt/4pt]black ChemLLM 12.83 0.869 10.34 0.979 0.851 0.849 0.773
gray!50[3pt/4pt]black ChemDFM 6.69 0.743 16.62 0.874 0.714 0.807 0.617
13.64 0.898 11.73 0.841 0.892 0.928 0.819
gray!50[3pt/4pt]black ChemCrow (+nach0) 30.52 0.976 11.91 0.990 0.947 0.977 0.905
14.60 0.962 16.10 0.981 0.891 0.942 0.805
LLM-based general models
InterLM2-Chat-7B 0.73 0.743 19.05 0.706 0.781 0.764 0.589
24.31 0.957 11.03 0.895 0.945 0.952 0.909
gray!50[3pt/4pt]black GPT-3.5 1.02 0.836 17.27 0.752 0.788 0.803 0.595
32.93 0.969 7.76 0.991 0.952 0.968 0.913
gray!50[3pt/4pt]black GPT-4o 0.30 0.839 17.68 0.751 0.743 0.646 0.493
32.85 0.970 7.72 0.990 0.953 0.970 0.914
gray!50[3pt/4pt]black Gemini-2.5-preview-05-20 4.19 0.790 17.20 0.864 0.765 0.720 0.572
39.02 0.986 6.63 0.961 0.973 0.989 0.944
gray!50[3pt/4pt]black Qwen2.5-7B-Instruct 0.02 0.670 30.80 0.668 0.522 0.396 0.289
37.55 0.971 7.64 0.992 0.961 0.973 0.922
Ours 70.28 0.998 2.27 0.997 0.990 0.996 0.989
Table 9: Comparison with LLM-based models on USPTO-50K. Space Setting indicates whether the model is provided with Top-4 reactant candidates and additional domain knowledge anchors (i.e., reaction types, reaction center) from specialized models. Since T5Chem and ChemFormer process the input by retaining only the product SMILES and directly predicting reactant SMILES, we do not report their performance with space. In contrast, ChemLLM relies on the four candidate reactants for selection and thus only its performance with space is reported.
The prompt for nach0 and ChemCrow.
System: You are a helpful AI Assistant that provides well-reasoned and detailed responses. You first think about the reasoning process as an internal monologue and then provide the user with the answer. Respond in the following format: $\n... \n</think>\n$$\n...\n</answer>$ ....................................................................................................................................................................... Assistant: You are an experienced chemist analyzing chemical retrosynthesis. Given the standard SMILES representation of the product is: {Product Standard SMILES}, Its automapping version is: {Product Mapped SMILES}, The IUPAC name of the product is: {IUPAC Name}. The possible reaction type for synthesizing this product is: {Reaction Type}. The possible reaction center for synthesizing this product is: {Reaction Center}. Reactants Candidate Set:
@@ 7 Details of Comparison Methods
The prompt for our Retro-Expert training.
@@
The prompt for ChemDFM.
You are an experienced chemist analyzing chemical retrosynthesis. Given the standard SMILES representation of the product is: {Product Standard SMILES}, Its automapping version is: {Product Mapped SMILES}, The IUPAC name of the product is: {IUPAC Name}. The possible reaction type for synthesizing this product is: {Reaction Type}. The possible reaction center for synthesizing this product is: {Reaction Center}. Reactants Candidate Set:
@@
The prompt for general LLM.
8 Details of Web Lab Experiments
8.1 Synthesis of 3-(2-ethoxyphenyl)thiophene
Figure 6: The synthesis path of 3-(2-ethoxyphenyl)thiophene.
(a) 1H NMR spectrum
(b) 13C NMR spectrum
8.2 Synthesis of 1-(4-ethoxyphenyl)ethanone
Figure 8: The synthesis path of 1-(4-ethoxyphenyl)ethanone.
9 Additional Ablation Study
This section presents additional ablation studies.
Reward hacking. Due to the probabilistic bias in the Varying-N candidate results provided by small models, i.e., the correct answer appears in the Top-1 position with significantly higher probability than in other positions. When using result correctness as the reward, the LLM will directly select the first candidate to continuously maximize its reward. To validate the effectiveness of our candidate position shuffling strategy, we visualized the positional distribution of correct answers in the small model’s N candidates and the distribution of positions selected by the LLM. As shown in Figure 10, using the original N candidates from small models results in severe reward hacking, whereas our position shuffling strategy effectively mitigates this issue.
Varying N Candidate. We analyze the impact of using different N values for candidate results during inference on LLM reasoning performance in Table 10. As N increases, the Top-1 reactant prediction accuracy gradually improves. When five or more candidates are provided, the model’s performance stabilizes at 70.8%. This indicates that the model can effectively perform critical reasoning based on the candidate set and generate the optimal answer. To ensure a fair comparison while maintaining competitive performance, we use the Top-4 candidate for reactant-related decision space, as chemical LLMs such as ChemLLM rely on four options for selection.
10 Results Analysis
@@
10.1 Successful Reasoning (Selection Scene)
We demonstrate the reasoning processes generated by our Retro-Expert across diverse reaction types (in Figure 11-16). This indicates Retro-Expert’s fundamental competency in executing chemically valid reasoning and making accurate selections within a high-quality chemical action space.
@@ 10.1 Successful Reasoning (Selection Scene)
N Candidates 1 3 4 5 7
Reactant Top-1 69.9 70.2 70.3 70.8 70.8
$\Delta$ +2.7 +3.0 +3.1 +3.6 +3.6
Table 10: The effect of the N value in small models’ N candidates on large model performance. $\Delta$ denotes the improvement relative to the Top-1 accuracy of small models’ reactant prediction (67.2%).
10.2 Successful Reasoning (Generation Scene)
Furthermore, we showcase Retro-Expert’s capacity for self-reflection and generation. Crucially, when confronted with suboptimal candidates from all specialized models, Retro-Expert engages in critical analysis and autonomously synthesizes a novel, chemically sound solution (in Figure 17-19).
@@ 10.3 Failure Analysis
(b) Failure case for multiple possible paths.
@@
11 Details of Interpretability Evaluation
@@
11.1 Details of Evaluation Metrics
@@
Mechanism Accuracy ($\mathbb{MA}$) System: You are an expert in chemical retrosynthetic analysis, responsible for evaluating the quality of reasoning processes generated by retrosynthetic models. #INSTRUCTIONS: Based on the following the Generated Reasoning Process, conduct a professional assessment across the follow dimensions: Mechanism Accuracy: Assess whether the reaction mechanism described in the reasoning aligns with established chemical principles. [Score 1]: Incorrect mechanism. [Score 2]: Partially correct but with significant deviations. [Score 3]: Partially correct with minor deviations. [Score 4]: Largely consistent. [Score 5]: Fully consistent. Evaluations must be grounded in given Authentic Chemical Synthesis Information and the results must be strictly formatted according to the specified Output Format. [Output Format Example] Mechanism Accuracy:2 ....................................................................................................................................................................... Assistant: Please evaluate the mechanism accuracy of the Generated Reasoning Process, based on the given Authentic Chemical Synthesis Information. {Generated Reasoning Process} Authentic Chemical Synthesis Information: The product’s standard SMILES representation is {Product Standard SMILES}; Its automapping version is {Product Mapped SMILES}; The IUPAC name of the product is {IUPAC Name}. The reaction type for synthesizing this product is: {The Golden Reaction Type}. The reaction center of the product molecule is: {The Golden Reaction Center}. The reactants for synthesizing the product molecule are: {The Golden Reactants}.
The prompt for Mechanism Accuracy ($\mathbb{MA}$).
@@
Factual Correctness ($\mathbb{FC}$) System: You are an expert in chemical retrosynthetic analysis, responsible for evaluating the quality of reasoning processes generated by retrosynthetic models. #INSTRUCTIONS: Based on the following the Generated Reasoning Process, conduct a professional assessment across the follow dimensions: Factual Correctness: Evaluate the presence of factual errors concerning chemical principles within the reasoning description. [Score 1]: Major factual errors. [Score 2]: Several errors. [Score 3]: minor error(s). [Score 4]: Essentially free of errors. [Score 5]: Completely accurate. Evaluations must be grounded in given Authentic Chemical Synthesis Information and the results must be strictly formatted according to the specified Output Format. [Output Format Example] Factual Correctness:2 ....................................................................................................................................................................... Assistant: Please evaluate the factual correctness of the Generated Reasoning Process, based on the given Authentic Chemical Synthesis Information. {Generated Reasoning Process} Authentic Chemical Synthesis Information: The product’s standard SMILES representation is {Product Standard SMILES}; Its automapping version is {Product Mapped SMILES}; The IUPAC name of the product is {IUPAC Name}. The reaction type for synthesizing this product is: {The Golden Reaction Type}. The reaction center of the product molecule is: {The Golden Reaction Center}. The reactants for synthesizing the product molecule are: {The Golden Reactants}.
The prompt for Factual Correctness ($\mathbb{FC}$).
@@
Logical Consistency ($\mathbb{LC}$) System: You are an expert in chemical retrosynthetic analysis, responsible for evaluating the quality of reasoning processes generated by retrosynthetic models. #INSTRUCTIONS: Based on the following the Generated Reasoning Process, conduct a professional assessment across the follow dimensions: Logical Consistency: Determine if conclusions follow logically from premises, analyses, or assumptions without contradictory leaps. [Score 1]: Conclusion contradicts premises/analysis. [Score 2]: Partially consistent but with significant logical leaps. [Score 3]: Partially consistent with only minor leaps. [Score 4]: Largely fluent and natural. [Score 5]: Completely fluent and natural. Evaluations must be grounded in given Authentic Chemical Synthesis Information and the results must be strictly formatted according to the specified Output Format. [Output Format Example] Logical Consistency:2 ....................................................................................................................................................................... Assistant: Please evaluate the logical consistency of the Generated Reasoning Process, based on the given Authentic Chemical Synthesis Information. {Generated Reasoning Process} Authentic Chemical Synthesis Information: The product’s standard SMILES representation is {Product Standard SMILES}; Its automapping version is {Product Mapped SMILES}; The IUPAC name of the product is {IUPAC Name}. The reaction type for synthesizing this product is: {The Golden Reaction Type}. The reaction center of the product molecule is: {The Golden Reaction Center}. The reactants for synthesizing the product molecule are: {The Golden Reactants}.
The prompt for Logical Consistency ($\mathbb{LC}$).
11.2 Details of Human Study
@@
12 Details of Out-of-Distribution Analysis
@@
The prompt for Out-of-Distribution Analysis.